# ATRNet-STAR: A Large Dataset and Benchmark Towards Remote Sensing Object Recognition in the Wild

### v3

@@ 2.2 Review of SAR Target Classification Algorithms
Early target classification relies on hand-crafting features and classifiers, while contemporary methodologies are predominantly anchored in deep learning-based methods. According to their different feature representations, we divided target recognition methods into four distinct paradigms: traditional features, deep learning, feature fusion, and classification with metadata. In an era increasingly dominated by data-driven algorithms, the availability of a rich, standardized, and usable dataset is becoming increasingly pivotal.
@@
Traditional features. Previous studies have proposed a wide range of valuable features, encompassing geometric structure features [54], electromagnetic-scattering features [55, 56, 57, 58], time-frequency features [59], local descriptors [7, 60], and sparse representations [61]. These features are particularly tailored to the special coherent imaging mechanisms. For instance, SAR-HOG [60] employs a ratio-based gradient computation rather than the traditional differential gradient approach, and Attributed Scattering Centers (ASC) provides a feature set including location, geometry, and polarization information. Notably, extracting electromagnetic-scattering features relies on complex images with phase information, underscoring the necessity of providing comprehensive data formats.
@@
Deep learning. Deep learning algorithms excel at uncovering intricate patterns within datasets, particularly when ample data is available. Nowadays, these algorithms achieve remarkable accuracy, reaching nearly 99% on MSTAR’s basic Standard Operating Condition (SOC) settings. However, the Extended Operating Condition (EOC), which includes robustness and few-shot, remains a significant challenge. Convolutional Neural Network (CNN) [62, 63, 64] is a common network in SAR target classification. All-convolutional network (A-ConvNet) [65] was proposed to reduce overfitting and had high accuracy under many experiment settings but is not robust to noise. In recent years, Transformer [66, 67, 68] emerged as a powerful alternative, offering new perspectives on feature extraction. After a decade of significant progress, SAR target recognition now incorporates many advanced deep learning techniques such as data augmentation [69, 70], attention mechanism [71, 72], domain alignment [53], reinforcement Learning [73, 74], transfer learning [75], meta-learning [76, 77, 78], physical deep learning [79, 80, 81], semi-supervised learning [82], and self-supervised learning [83, 29]. The majority of research endeavors are based on the MSTAR dataset, and a subset of studies also create the different non-open source settings of OpenSARShip and FUSAR-Ship datasets. Nonetheless, the reliance on outdated datasets and the lack of standardized benchmarks are impeding the advancement of SAR target recognition in the 2020s.
@@
Feature fusion. In order to improve the robustness and interpretability of deep learning methods, many feature fusion methods combining traditional and deep features have been proposed. Kechagias et al. [84] discussed the fusion of deep learning and sparse coding. FEC [85] used deep CNN combined with ASC features derived from complex data, incorporating both amplitude and phase information. Li et al. [86] combined scattering centers (SC) with graph convolutional network. Besides feature fusion methods, ESF [87] applied ASC features to guide deep model training. EFTL [88] proposed using the ASC model as weight initialization of a complex convolutional network. PIHA [89] proposed a hybrid attention mechanism with ASCs and the deep learning framework. LDSF [90] constructed a heterogeneous graph for the first time to fully exploit both the scattering information of the target components and their interactions. However, many SAR target datasets only provide magnitude images, ignoring phase information and quantization issues.
@@
Classification with metadata. Imaging angles and target-related metadata are the two most prevalent categories of ancillary data. MSTAR and SAMPLE datasets furnish the imaging angles for each respective image. RotANet [91] combined predicting the rotational pattern of an image sequence with self-supervised learning. Zhang et al. [92] investigated the problem of optimal image imaging angle selection in few-shot recognition based on the MSTAR dataset. AaDRL [73] explored the SAR active recognition framework to fully exploit target characteristic variations at different azimuth angles based on the SAMPLE dataset. In addition, multi-view methods [93] were proposed with image sequences from different imaging angles. Since the target types are known, the researchers used 3D models and shape parameters as metadata for fine-grained tasks. For example, SARbake [94] used 3D CAD models and projection relations for semantic segmentation, and DBAE [95] utilized size and structure attributes for zero-shot classification. Consequently, a SAR target dataset enriched with pertinent metadata, encompassing both imaging parameters and detailed target characteristics, becomes an indispensable step in pursuing accurate and robust SAR target recognition.
@@
Experimental settings. Indeed, the MSTAR dataset has more than 50 types of experimental configurations for SOC and EOC, as detailed in [11]. These configurations include image angle variation, target version/configure variation, noise corruption, occlusion, resolution variation, and background clutter variation [11, 53]. In a similar vein, other SAR classification datasets, such as OpenSARShip and FUSAR-Ship, lack a standard experimental benchmark, and researchers also devise different bespoke settings that leverage subsets of the target class for classification [81].
@@
Evaluation metrics. Classification accuracy and confusion matrix are the primary metrics employed for evaluating the performance of classification models. Accuracy quantifies the proportion of correctly classified samples relative to the total number of samples in the dataset. The confusion matrix provides a detailed correspondence between the predicted and true class labels, with each column representing the predicted class and each row representing the true class. Besides, the Receiver Operating Characteristic (ROC) curve is a valuable tool for assessing used for binary classification and detection tasks such as suppressing background clutter and outlier rejection [65].
@@
We noticed a pronounced preference for the MSTAR dataset among SAR classification algorithms. Most algorithms are based on the MSTAR dataset, with a small number using other datasets such as SAMPLE, OpenSARShip, and FUSAR-Ship. This preference is attributed to the distinctive benefits of the MSTAR dataset released in the 1990s. As depicted in Fig. 4 (a), studies requiring complex data have opted for the MSTAR dataset, whereas other datasets that only provide magnitude data cannot be used to validate their algorithms. In addition, information about vehicle types in the dataset is also available on the web. Compared to other datasets with long-tailed distributions, the MSTAR dataset has a more balanced target sample and diverse imaging conditions. Therefore, MSTAR has accumulated a large number of algorithms and experimental settings. However, this dataset has not released all target samples, and the extant algorithms have achieved peak performance in common popular settings [11], such as classification with ten target types, imaging angle variation, and version variation. The ideal acquisition conditions of the MSTAR dataset cannot show that the recognition problem is truly well solved. Scholars aim to address these limitations by proposing novel simulation experimental settings. However, as shown in Fig. 4 (b), some simulation experimental settings may not accurately reflect the real-world conditions. Besides, ensuring sample diversity is fraught with challenges due to the high data collection and labeling costs. These limitations collectively impose significant constraints on the advancement of SAR ATR techniques. Constructing a new open-source dataset benchmark with diverse target categories and imaging conditions is a challenging cornerstone of excellent and stable SAR ATR techniques.
@@ 3 Building ATRNet-STAR: Why, What, How
In this section, we present ATRNet-STAR from three perspectives: (1) Why we constructed this dataset in the 2020s; (2) What we did to ensure its diversity and standardization; (3) How the diversity and statistical analysis of this dataset to support SAR ATR.
@@
The emergence of new datasets and papers has led to rapid progress in various areas of SAR ATR, demonstrating the attention given to this issue in recent years. However, a large-scale dataset with various target categories and imaging conditions is still lacking because of SAR’s expensive costs and annotation. We aim to achieve as much diversity as possible in target categories, scene settings, and sensor conditions. Such multidimensional design establishes it as an instrumental resource for advancing SAR ATR research. Its principal values for the SAR ATR are fourfold:
@@ 3.1 Construct program
Target. First of all, according to the Chinese and European vehicle classification standards [96, 97], we confirm the main target classes based on uses and structures of vehicles, including people carriers (car and bus), goods carriers (truck), and special purpose vehicles (special). Then, we subdivide each class into various subclasses based on size and special structure and collect as many different types as possible. Finally, ATRNet-STAR has 4 classes, 21 subclasses, and 40 types. It covers most civilian vehicle categories and has the most richness on civil target classes compared to other SAR vehicle datasets.
Scene. Most previous target datasets were collected in open scenes such as grasslands, focusing on target characteristics and lacking significant background interference. Therefore, we discuss both difficult and simple scenes222Since we mainly consider single target characteristics, discussing densely distributed targets such as car parks are not included in this version. in Fig. 10. The urban and factory have varying occlusion degrees and other strong scattering targets, such as buildings and roadside trees. Dense trees in woodlands can weaken target scattering or increase the intensity of non-target scattering. Bare soil and sandstone are clean backgrounds, although sandstone occasionally has strong scattering background points. These interferences have a stronger effect on the target features and are more compatible with real-world challenging recognition tasks.
Sensor. The common sensor conditions are variations in depression and azimuth angle, and we collect target samples at different angles as shown in Fig. 11. We provide a balanced sample under different depression angles, including all target types, compared to MSTAR’s several targets. However, our azimuth angles are not as dense as MSTAR in most scenes due to the limitation of strimap imaging. In addition, we provide data in different polarizations and bands to study their effects on SAR target characteristics. We simultaneously gather data from satellites and airborne to study the platforms and resolutions.
@@ 3.2 Data acquisition
Imaging. We use the Unmanned Aerial Vehicle (UAV) platform to carry sensor equipment and collect data. Two antennas acquire quad polarization radar echo in the X and Ku bands. Besides, the Position and Orientation System (POS) device provides the Global Positioning System (GPS) and Inertial Measurement Unit (IMU) information for motion compensation. After imaging processing, we have slant range complex images. Ground range data is obtained by ground projection based on slant range amplitude images and POS information.
Annotation. Target categories and coordinates are annotated using rectangular box labels based on optical reference images and deployment records. Focusing on individual target signatures, all objects maintain specified separation distances during placement. Besides our vehicles, other vehicles in the scene are labeled as “other”. After labeling, we acquired the target slices and added a random offset.
Product. We offer data products in two coordinate systems. The distance dimension of the slant range data is the line of sight direction, and we provide original complex data. The ground range images is projected to the ground truth distance and is processed with nonlinear quantization. The corresponding annotation files include basic image information and target, scene, and sensor parameters.
@@ 3.3 Professional system
We construct a target taxonomic system to facilitate subsequent extensions and integration with other datasets. Data format and metadata are also discussed, considering the specificity of SAR imaging regarding view, steps, and operating conditions.
@@
Taxonomic system. A taxonomic system is crucial when creating a large-scale fine-grained dataset [99, 26, 98]. Motivated by MSTAR’s military taxonomic system [98] and FUSAR-Ship’s ship taxonomic system [26], our hierarchical taxonomic system (class subclass type) are based on vehicle’s purpose, structure, size and mass referred to Chinese and European vehicle classification standards [96, 97]. As shown in Fig. 5 (a), we first create 4 class based on the vehicle’s different purposes, such as carrying passengers (car and bus), cargo (truck), and special purposes (special), and passenger and cargo dual-purpose pickup is classified into the truck. Then, the system is further expanded into 21 subclasses through differentiation of structural, size, and mass specifications. For example, cars are based on size and passenger capacity, and trucks depend on length, mass, and special structure. Finally, we collect the 40 vehicle types (e.g., EV160B and QQ3 types from different producers and brands in the mini car class.) with size information in Fig. 7. The target classes of SAR vehicle datasets are summarized in Fig. 5 (b). This hierarchical architecture enhances standardization in data acquisition protocols and supports modular dataset expansion.
@@
Data format. Due to the unique imaging mechanisms, processing steps, and image properties of SAR, we offer different levels of products to support the comprehensive study of the impact of imaging steps on recognition problems. We provide magnitude images in the ground range coordinate system with 8-bit and 32-bit uint formats. The magnitude images with 16-bit uint and complex data with 32-bit float in the slant range coordinate system are also available. In subsequent experimental benchmarks, we use 8-bit ground distance magnitude images and 32-bit slant distance complex data. We also perform other processing based on the slant distance data, such as phase gradient autofocus and polarization decomposition, but the current version of this dataset does not include the processed data after these steps. Raw echoes are also stored, but due to the large storage (25TB), they will not be available for upload and download. Researchers can contact us to obtain raw echo if needed. We hope these products will help the researchers easily work with our dataset while providing more tractability to explore and discuss specific imaging mechanisms and image properties for SAR recognition applications.
@@
Metadata system. Considering the effects of target, scene, and sensor on SAR images, we add many factors to the annotation to facilitate various recognition methods with metadata. Image information includes basic image sizes and formats. The target attributes include taxonomic systems, target size, and image position. Although we discuss single target characteristics with classification tasks, rectangular boxes also support detection tasks. Sensor parameters have the depression angle, target azimuth angle, band, polarization, and resolution in two dimensions. However, we hope that researchers consider the difficulty and accuracy of obtaining metadata when making full use of them. For example, there is the potential for angular confusion when measuring target azimuths and the potential for vehicle size to change after modifications.

### v4

@@ 2.2 Review of SAR Target Classification Algorithms
Early target classification relies on hand-crafting features, while contemporary methodologies are predominantly anchored in deep learning-based methods. According to their different feature representations, we divided target recognition methods into four distinct paradigms: traditional features, deep learning, feature fusion, and classification with metadata. In an era increasingly dominated by data-driven algorithms, the availability of a rich, standardized, and usable dataset is becoming increasingly crucial.
@@
Traditional features - Previous studies have proposed a wide range of valuable features, encompassing geometric structure features [55], electromagnetic-scattering features [56, 57, 58, 59], time-frequency features [60], local descriptors [7, 61], and sparse representations [62]. These features are particularly tailored to the special coherent imaging mechanisms. For instance, SAR-HOG [61] employs a ratio-based gradient computation rather than the traditional differential gradient approach, and Attributed Scattering Centers (ASC) provides a feature set including location, geometry, and polarization information. Notably, extracting electromagnetic-scattering features relies on complex images with phase information, underscoring the necessity of providing comprehensive data formats.
@@
Deep learning - Deep learning algorithms can uncover intricate patterns within datasets, particularly when ample data is available. Nowadays, these algorithms achieve remarkable accuracy, reaching nearly 99% on MSTAR’s basic Standard Operating Condition (SOC) settings. However, the Extended Operating Condition (EOC), which includes robustness and few-shot, remains a significant challenge. Convolutional Neural Network (CNN) [63, 64, 65] is a common network in SAR target classification. All-convolutional network (A-ConvNet) [66] was proposed to reduce overfitting and had high accuracy under many experiment settings but is not robust to noise. In recent years, Transformer [67, 68, 69] emerged as a powerful alternative, offering new perspectives on feature extraction. After a decade of significant progress, SAR target recognition now incorporates many advanced deep learning techniques such as data augmentation [70, 71], attention mechanism [72, 73], domain alignment [54], reinforcement Learning [74, 75], transfer learning [76], meta-learning [77, 78, 79], physical deep learning [80, 81, 82], semi-supervised learning [83], and self-supervised learning [84, 30]. The majority of research endeavors are based on the MSTAR dataset, and a subset of studies also create the different non-open source settings of OpenSARShip and FUSAR-Ship datasets. Nonetheless, the reliance on outdated datasets and the lack of standardized benchmarks are impeding the advancement of SAR target recognition in the 2020s.
@@
Feature fusion - In order to improve the robustness and interpretability of deep learning methods, many feature fusion methods have been proposed combining traditional and deep features. Kechagias et al. [85] discussed the fusion of deep learning and sparse coding. FEC [86] used deep CNN combined with ASC features derived from complex data, incorporating both amplitude and phase information. Li et al. [87] combined scattering centers (SC) with graph convolutional network. Besides feature fusion methods, ESF [88] applied ASC features to guide deep model training. EFTL [89] proposed using the ASC model as weight initialization of a complex convolutional network. PIHA [90] proposed a hybrid attention mechanism with ASCs and the deep learning framework. LDSF [91] constructed a heterogeneous graph for the first time to fully exploit both the scattering information of the target components and their interactions. However, many SAR target datasets only provide magnitude images, ignoring phase information and quantization issues.
@@
Classification with metadata - Imaging angles and target-related metadata are the two most prevalent categories of ancillary data. MSTAR and SAMPLE datasets furnish the imaging angles for each respective image. RotANet [92] combined predicting the rotational pattern of an image sequence with self-supervised learning. Zhang et al. [93] investigated the problem of optimal image imaging angle selection in few-shot recognition based on the MSTAR dataset. AaDRL [74] explored the SAR active recognition framework to fully exploit target characteristic variations at different azimuth angles based on the SAMPLE dataset. In addition, multi-view methods [94] were proposed with image sequences from different imaging angles. Since the target types are known, the researchers used 3D models and shape parameters as metadata for fine-grained tasks. For example, SARbake [95] used 3D CAD models and projection relations for semantic segmentation, and DBAE [96] utilized size and structure attributes for zero-shot classification. Consequently, a SAR target dataset enriched with pertinent metadata, encompassing both imaging parameters and detailed target characteristics, becomes an indispensable step in pursuing accurate and robust SAR target recognition.
@@
Experimental settings - Indeed, the MSTAR dataset has more than 50 types of experimental configurations for SOC and EOC, as detailed in [11]. These configurations include image angle variation, target version/configure variation, noise corruption, occlusion, resolution variation, and background clutter variation [11, 54]. In a similar vein, other SAR classification datasets, such as OpenSARShip and FUSAR-Ship, lack a standard experimental benchmark, and researchers also devise different bespoke settings that leverage subsets of the target class for classification [82].
@@
Evaluation metrics - Classification accuracy and confusion matrix are the primary metrics employed for evaluating the performance of classification models. Accuracy quantifies the proportion of correctly classified samples relative to the total number of samples in the dataset. The confusion matrix provides a detailed correspondence between the predicted and true class labels, with each column representing the predicted class and each row representing the true class. Besides, the Receiver Operating Characteristic (ROC) curve is a valuable tool for assessing used for binary classification and detection tasks such as suppressing background clutter and outlier rejection [66].
@@
We noticed a pronounced preference for the MSTAR dataset among SAR classification algorithms. Most algorithms are based on the MSTAR dataset, with others using datasets such as SAMPLE, OpenSARShip, and FUSAR-Ship. This preference is attributed to the distinctive benefits of the MSTAR dataset released in the 1990s. As depicted in Fig. 4 (a), studies requiring complex data have opted for the MSTAR dataset, whereas other datasets that only provide magnitude data cannot be used to validate their algorithms. In addition, information about vehicle types in the dataset is also available on the web. The MSTAR dataset has a more balanced target sample and diverse imaging conditions compared to other datasets with long-tailed distributions. Therefore, MSTAR has accumulated a large number of algorithms and experimental settings. However, this dataset has not released all target samples, and many algorithms have achieved peak performance in common popular settings [11], such as classification with ten target types, imaging angle variation, and version variation. The ideal acquisition conditions of the MSTAR dataset cannot show that the recognition problem is truly well solved. Scholars aim to address these limitations by proposing novel simulation experimental settings. However, as shown in Fig. 4 (b), some simulation experimental settings may not accurately reflect the real-world conditions. Besides, ensuring sample diversity is challenging due to the high data collection and labeling costs. These limitations collectively impose significant constraints on the advancement of SAR ATR techniques. Constructing a new open-source dataset benchmark with diverse target classes and imaging conditions is an urgent cornerstone of excellent and stable SAR ATR techniques.
@@ 3 Building ATRNet-STAR: Why, What, How
In this section, we present and discuss ATRNet-STAR from three perspectives: (1) Why we constructed this dataset in the 2020s; (2) What we did to ensure its diversity and standardization; (3) How the diversity and statistical analysis of this dataset to support SAR ATR.
@@
The emergence of new datasets and papers has led to rapid progress in various areas of SAR ATR, demonstrating the attention given to this issue in recent years. However, a large-scale dataset with various target classes and imaging conditions is still lacking because of SAR’s expensive costs and annotation. We aim to achieve as much diversity as possible in target classes, scene settings, and sensor conditions. Such multidimensional design establishes it as an instrumental resource for advancing SAR ATR research. Its principal values for the SAR ATR are fourfold:
@@ 3.1 Construct program
Target - First of all, according to the Chinese and European vehicle classification standards [97, 98], we confirm the main target classes based on uses and structures of vehicles, including people carriers (car and bus), goods carriers (truck), and special purpose vehicles (special). Then, we subdivide each class into various subclasses based on size and special structure and collect as many different types as possible. Finally, ATRNet-STAR has 4 classes, 21 subclasses, and 40 types. It covers most civilian vehicle categories and has the most richness on civil target classes compared to other SAR vehicle datasets.
Scene - Most previous target datasets were collected in open scenes such as grasslands, focusing on target characteristics and lacking significant background interference. Therefore, we discuss both difficult and simple scenes222Since we mainly consider single target characteristics, discussing densely distributed targets such as car parks are not included in this version. in Fig. 10. The urban and factory have varying occlusion degrees and other strong scattering targets, such as buildings and roadside trees. Dense trees in woodlands can weaken target scattering or increase the intensity of non-target scattering. Bare soil and sandstone are clean backgrounds, although sandstone occasionally has strong scattering background points. These interferences have a stronger effect on the target features and are more compatible with real-world challenging recognition tasks.
Sensor - The common sensor conditions are variations in depression and azimuth angle, and we collect target samples at different angles as shown in Fig. 11. We provide a balanced sample under different depression angles, including all target types, compared to MSTAR. However, our azimuth angles are not as dense as MSTAR in most scenes due to the limitation of strimap imaging. In addition, we provide data in different polarizations and bands to study their effects on SAR target characteristics. We simultaneously gather some data from satellites to study the platforms and resolutions.
@@ 3.2 Data acquisition
Imaging - We use the Unmanned Aerial Vehicle (UAV) platform to carry sensor equipment and collect data. Two antennas acquire quad polarization radar echo in the X and Ku bands. Besides, the Position and Orientation System (POS) device provides the Global Positioning System (GPS) and Inertial Measurement Unit (IMU) information for motion compensation. After imaging processing, we have slant range complex images. Ground range data is obtained by ground projection based on slant range amplitude images and POS information.
Annotation - Target classes and coordinates are annotated using rectangular box labels based on optical reference images and deployment records. Because of focusing on individual target signatures, all objects maintain specified separation distances during placement. Besides our vehicles, other vehicles in the scene are labeled as “other”. After labeling, we acquired the target slices and added a random offset.
Product - We offer data products in two coordinate systems. The distance dimension of the slant range data is the line of sight direction, and we provide original complex data. The ground range images is projected to the ground truth distance and is processed with nonlinear quantization. The corresponding annotation files include basic image information and target, scene, and sensor parameters.
@@ 3.3 Professional system
We construct a target taxonomic system to facilitate subsequent extensions and integration with other datasets. Data format and annotation metadata are also discussed, considering the specificity of SAR imaging.
@@
Taxonomic system - A taxonomic system is crucial when creating a large-scale fine-grained dataset [100, 27, 99]. Motivated by MSTAR’s military taxonomic system [99] and FUSAR-Ship’s ship taxonomic system [27], our hierarchical taxonomic system (class subclass type) are based on vehicle’s purpose, structure, size and mass referred to Chinese and European vehicle classification standards [97, 98]. As shown in Fig. 5 (a), we first create 4 classes based on the vehicle’s different purposes, such as carrying passengers (car and bus), cargo (truck) and special purposes (special), and dual-purpose pickup is classified into the truck. Then, the system is further expanded into 21 subclasses according to structural, size, and mass specifications. For example, cars are based on size and passenger capacity, and trucks depend on length, mass, and special structure. Finally, we collect the 40 vehicles types (e.g., EV160B and QQ3 types from different producers and brands in the mini car class.) with size information in Fig. 7. The target classes of SAR vehicle datasets are summarized in Fig. 5 (b). This hierarchical architecture enhances standardization in data acquisition protocols and supports modular dataset expansion.
@@
Data format - Due to the unique imaging mechanisms, processing steps, and image properties of SAR, we offer different levels of products to support the comprehensive study of the impact of imaging steps on recognition problems. We provide magnitude images in the ground range coordinate system with 8-bit and 32-bit uint formats. The magnitude images with 16-bit uint and complex data with 32-bit float in the slant range coordinate system are also available. In subsequent experimental benchmarks, we use 8-bit ground distance magnitude images and 32-bit slant distance complex data. We also perform other processing based on the slant distance data, such as phase gradient autofocus and polarization decomposition, but the current version of this dataset does not include the processed data after these steps. Raw echoes are also stored, but due to the large storage (25TB), they will not be available for upload and download. Researchers can contact us to obtain raw echo if needed. We hope these products will help the researchers easily work with our dataset while providing more tractability to explore and discuss specific imaging mechanisms and image properties for SAR recognition applications.
@@
Metadata system - Considering the effects of target, scene, and sensor on SAR images, we add many factors to the annotation to facilitate various recognition methods with metadata. Image information includes basic image sizes and formats. The target attributes include taxonomic systems, target size, and image position. Although we discuss single target characteristics with classification tasks, rectangular boxes also support detection tasks. Sensor parameters have the depression angle, target azimuth angle, band, polarization, and resolution in two dimensions. However, we hope that researchers consider the difficulty and accuracy of obtaining metadata when making full use of them. For example, there is the potential for angular confusion when measuring target azimuths and the potential for vehicle size to change after modifications.